OpenStudy (anonymous):

2x-3 | >=5 *sn there is a line under > this symbol

OpenStudy (amistre64):

2x -3 >= 5 and -(2x -3) >= 5 2x >=8 -2x >= 2 x >=4 x<= -1 solution is : (-inf,-1] or [4,inf)

OpenStudy (anonymous):

x >=4 or x<= -1 can i write it like this ?

OpenStudy (amistre64):

yes, thats good if the 'program' accepts it :)

OpenStudy (anonymous):

okay thanks

OpenStudy (amistre64):

i wonder if I make a long equation for this answer if I get a badge lol\[25 \int\limits_{} \frac{x-2}{a^2+x^2}dx \rightarrow F(x)\]